# Cislunar Constellation Design for Space Situational Awareness with Time-Expanded Facility Location Problem

### v1

@@ 4.2 Problem Formulation
The TE-$p$-MP aims to maximize the aggregated coverage of all targets across all time steps. We collect the visibility boolean $\bar{m}^{b}(\boldsymbol{r}_{j}(t),\boldsymbol{r}_{k}(t),i)$ from equation (9) corresponding to observer location $j$, pointing direction $i$, target $k$, at time $t$ into a visibility tensor $M^{b}\in\mathbb{B}^{m\times n\times\ell\times q}$,
@@
The objective (13a) has two terms; the primary contribution comes from the first term, which maximizes the fraction of visible targets at each time-step, and takes its summation. Note that this is different from simply taking $\sum_{t=1}^{\ell}\sum_{k=1}^{q}\sum_{i=1}^{m}\sum_{j=1}^{n}M^{b}_{ijtk}X_{ijt}$, which would allow the optimizer to capitalize on observing the same target by multiple observers rather than observing as many targets as possible. The second term is a discount factor that penalizes the use of “expensive” observer locations, where $f\in\mathbb{R}^{n}$ is the vector of costs associated with making use of a given facility location, given by
@@
where $\nu_{j}$ is the stability index of the LPO, as defined by equation (4). The definition of $f_{j}$ with equation (14) scales the stability index, which can take values with two orders of magnitude of difference, to values around 1. Since prohibitively expensive candidate observer locations should be excluded from the problem to begin with, it is assumed that all candidate slots are at least acceptable to be part of the solution; as such, this term is simply meant to favor combinations of facilities that are more stable in the case that multiple combinations of facilities yield the same value from the first term.
@@
where constraint (15b) ensures this modified problem (15) to have the same optimal solution as the original problem (13). This is now a standard mixed-binary linear program (MBLP), which is amenable to standard solvers such as Gurobi [43] or CPLEX [44]. While the slack variable $\theta_{tk}$ is a continuous variable in $[0,1]$, the optimal solution to (15) has $\theta_{tk}=1$ for time $t$ and target $k$ if at least one observer is viewing $k$ at this time, and $\theta_{tk}=0$ if $\sum_{i=1}^{m}\sum_{j=1}^{n}M^{b}_{ijtk}X_{ijt}=0$. Thus, this slack variable may be interpreted as a binary variable, such that
@@
For convenience in comparing obtained solutions later, we define the fraction of observation demand that has been met across all time steps, denoted as $\Theta$, and given by
@@
The formulation (15) enjoys a few favorable properties; firstly, even though the slack variables $\theta$ increase the decision space, it is a continuous variable, mending itself well to standard linear program techniques. Secondly, despite the introduction of the time index $t$ and the decomposition of allocation into pointing directions $i$ and targets $k$, the general structure of the traditional $p$-median problem is preserved; this enables efficient use of Lagrangean relaxation, which will be introduced in Section 5.
5 Lagrangean Method for the Time-Expanded $p$-Median Problems
One popular approach to solving complex binary linear programs (BLP) is the Lagrangean method (LM), which iteratively uses Lagrangean relaxation (LR) and problem-specific heuristics to seek the optimal solution by computing upper and lower bound solutions. The fundamental intuition behind LM is to apply LR to remove complicating constraints by adjoining them using Lagrange multipliers into the objective, in such a way that the original BLP may be decoupled into smaller subproblems that are easier to solve. Since we are dealing with a maximization problem, the solution to the relaxed problem provides an upper bound of the optimal solution sought; using the solution to the LR, problem-specific heuristics are employed to generate a feasible solution, corresponding to a feasible lower bound. This process is iterated, each time updating the Lagrange multipliers towards optimality, for example using a subgradient approach.
@@ 5 Lagrangean Method for the Time-Expanded $p$-Median Problems
For the $p$-Median problem, a popular LR is to relax constraints that pertain to allocations between $i$ and $j$, resulting in decoupled subproblems corresponding to each candidate facility location $j$; LR is particularly advantageous for the $p$-Median problem as these subproblems may be solved analytically. In this work, despite the introduction of the time indices $t$ and slack variables $\theta_{tk}$, we also arrive at LR expressions that achieve decomposition into subproblems corresponding to each $j$. We further provide a specialized set of heuristics to generate a competitive feasible lower bound solution.
@@
In the remainder of this Section, we begin with Section 5.1 where we discuss the LR formulation to the original problems, along with the approach to decompose the resulting LR problem into facility-wise subproblems and to solve each of these subproblems. This is followed by Section 5.2 which describes the heuristic scheme devised to generate a feasible solution based on the LR solution at each iteration. An additional set of heuristics for performing a neighborhood search in an attempt to improve this feasible solution is described in Section 5.3. In Section 5.4, the subgradient method used to update the multipliers is introduced. The processes over one iteration of the LM are summarized in Figure 8. Finally, in Section 5.5, we provide implementation considerations for implementing the proposed LM.
Solve relaxed problem to obtain upper bound (Section 5.1)Convert allocation to obtain lower bound (Section 5.2)Perform heuristic neighborhood search (Section 5.3)Update Lagrange multipliers (Section 5.4)
@@
Figure 8: Overview of a single iteration of Lagrangean method for the time-expanded $p$-Median problem
5.1 Upper Bound Solution via Lagrangean Relaxation
We specifically present the Lagrangean relaxation of the aggregated coverage problem. Relaxing constraints (13b) and (15b) with multipliers $\lambda_{jt}$ and $\eta_{tk}$ respectively, the Lagrangean relaxed problem is given by
$\max_{X,Y,\theta}\quad$ $\sum_{t=1}^{\ell}\sum_{k=1}^{q}\theta_{tk}-\dfrac{1}{\ell}\sum_{j=1}^{n}f_{j}Y_{j}-\sum_{j=1}^{n}\sum_{t=1}^{\ell}\lambda_{jt}\left(\sum_{i=1}^{m}X_{ijt}-1\right)-\sum_{t=1}^{\ell}\sum_{k=1}^{q}\eta_{tk}\left(\theta_{tk}-\sum_{i=1}^{m}\sum_{j=1}^{n}M^{b}_{ijtk}X_{ijt}\right)$ (18a)
s.t. $\lambda_{it},\eta_{tk}\geq 0\quad\forall t,j,k$ (18b)
(13c) -(13e), (15c)
@@ 5.1 Upper Bound Solution via Lagrangean Relaxation
The sought optimal decision variables to the relaxed problem (18), denoted by $\bar{X}$, $\bar{Y}$, and $\bar{\theta}$ generates an upper bound to problem (15).
@@ 5.1.1 Decomposition into Subproblems
We can further recognize that this relaxation decouples the problem on a per-candidate observer location basis, where for each candidate location $j$, the subproblem is given by
@@
$\lambda_{it},\eta_{tk}\geq 0\quad\forall t,k$ (21d)
Note that this subproblem is defined for each $j$, and has as variables $X_{ijt}\in\mathbb{B}^{m\times 1\times\ell}$, $Y_{j}\in\mathbb{B}$, and $\theta\in\mathbb{R}^{\ell\times k}$. We denote the decision on $X$ and $Y_{j}$ made with the $j^{\rm th}$ subproblem as $\bar{X}^{j}$ and $\bar{Y}_{j}^{j}$, which denote the allocations of the $j^{\rm th}$ observer to pointing directions over time, as well as whether the $j^{\rm th}$ observer should be used, respectively. We consider separately the cases where $\bar{Y}^{j}_{j}=0$ or $\bar{Y}^{j}_{j}=1$; the decision on whether $\bar{Y}^{j}_{j}=0$ or $\bar{Y}^{j}_{j}=1$ should be chosen is not made at the subproblem level, but is rather left for later when the LR solution is constructed by assembling solutions from each subproblem. If $\bar{Y}^{j}_{j}=0$, constraint (21b) trivially forces $\bar{X}^{j}_{ijt}=0$ $\forall i,t$. If $\bar{Y}^{j}_{j}=1$, the choice of each $\bar{X}^{j}_{ijt}$ $\forall i,t$ is driven by its coefficient in the objective, and thus the optimal value $\bar{X}_{ijt}^{j}$ is given by
@@
The corresponding objective of the $j^{\mathrm{th}}$ subproblem (21a) assuming $\bar{Y}_{j}=1$, denoted as $Z^{j}_{\mathrm{LR}}$, is given by
$Z^{j}_{\mathrm{LR}}=\sum_{i=1}^{m}\sum_{t=1}^{\ell}\left(\sum_{k=1}^{q}\eta_{tk}M^{b}_{ijtk}-\lambda_{jt}\right)\bar{X}_{ijt}^{j}-\dfrac{f_{j}}{\ell}+\sum_{t=1}^{\ell}\lambda_{jt}$ (23)
5.1.2 Construction of Lagrangean Relaxed Solution
To construct the LR solution, we first compute $Z^{j}_{\mathrm{LR}}$ for all $j=1,\ldots,n$ subproblems (21a) according to equation (23). We then define the permutation $\pi$ for the $n$ candidate observer locations such that
$Z^{\pi(1)}_{\mathrm{LR}}\geq Z^{\pi(2)}_{\mathrm{LR}}\geq\ldots\geq Z^{\pi(m)}_{\mathrm{LR}}$ (24)
Then, to enforce the $p$-median constraint (13c), variables $\bar{Y}$ are chosen such that the $p$ best facilities are chosen
@@ 5.1.2 Construction of Lagrangean Relaxed Solution
Finally, the upper bound objective of the Lagrangean relaxed problem (18a) $Z_{\mathrm{LR}}$ is given by substituting the decision variables to equation (15a),
$Z_{\mathrm{LR}}=\sum_{t=1}^{\ell}\sum_{k=1}^{q}\left(1-\eta_{tk}\right)\bar{\theta}_{tk}+\sum_{s=1}^{p}Z^{\pi(s)}_{\mathrm{LR}}$ (27)
5.2 Lagrangean Heuristics Reallocation
If the Lagrangean relaxed solution $(\bar{X},\bar{Y},\bar{\theta})$ is feasible, then it is the optimal solution $(X^{*},Y^{*},\theta^{*})$ to problem (15). This is unlikely as constraints (13b) and (15b) have been relaxed. Thus, we seek to construct a feasible lower bound solution $(\underline{X},\underline{Y},\underline{\theta})$ using the relaxed solution. The overall procedure is summarized in Algorithm 1.
@@ 5.2 Lagrangean Heuristics Reallocation
and we construct the set of all targets that are unobserved at this time, $\mathcal{K}^{0}_{t}$
@@
Equivalently, this set may be understood as the set of $j$ that still requires allocation to an appropriate pointing direction $i$. The next step consists of adding back allocations for $j\in\mathcal{J}_{t}^{0}$ for each time-step $t$. This corresponds to the GreedyAllocation() functions call in line 14 of algorithm 1, and is described in further detail in the next subsection. Finally, we fix the solution to satisfy constraint (15b) by iterating through $t$ and setting $\underline{\theta}_{tk}=0$ if
@@
Finally, the lower bound feasible solution $Z_{\mathrm{LH}}$ is given by
$Z_{\mathrm{LH}}=\sum_{t=1}^{\ell}\sum_{k=1}^{q}\left(1-\eta_{tk}\right)\underline{\theta}_{tk}+\sum_{j=1}^{n}f_{j}\underline{Y}_{j}$ (33)
@@ 5.2.1 Greedy Allocation
In a greedy process, the facility $j^{\star}$ with pointing direction $i^{\star}$ that results in the largest additional number of targets observed is selected via the expression
@@
This process is repeated, each time setting $X_{i^{\star}j^{\star}t}=1$ and removing $j^{\star}$ from $\mathcal{J}_{t}^{0}$, until $\mathcal{J}_{t}^{0}=\varnothing$. Algorithm 2 gives the pseudo-code for implementing this greedy allocation strategy.
@@
5.2.2 Full-Factorial Allocation Strategy
Typically, the number of observers $p$ in the context of the CSSA problem is relatively small, on the orders of 1s to 10s. As such, it may be computationally reasonable to spend more time sequentially allocating remaining observers $j\in\mathcal{J}^{0}_{t}$ to remaining targets $\mathcal{K}^{0}_{t}$. We refer to this process as the full-factorial allocation; Algorithm 3 shows the pseudo-code for its procedure. The while-loop from the greedy algorithm 2 is now replaced by the outer for-loop in line 2, which iterates through all possible permutations of indices of the set $\mathcal{J}^{0}_{t}$, denoted by $\Pi_{\mathcal{J}}$. For each permutation $\pi_{\mathcal{J}}\in\Pi_{\mathcal{J}}$, we go through indices $j$ in the order of $\pi_{\mathcal{J}}$, each time allocating $j$ to pointing direction $i$ in a greedy manner, such that

### v2

@@ 4.2 Problem Formulation
We formulate a problem to design a CSSA constellation that maximizes the aggregated coverage of all targets across all time steps. We collect the visibility boolean $\bar{m}^{b}(\boldsymbol{r}_{j}(t),\boldsymbol{r}_{k}(t),i)$ from equation (19) corresponding to observer location $j$, pointing direction $i$, target $k$, at time $t$ into a visibility tensor $M^{b}\in\mathbb{B}^{m\times n\times\ell\times q}$,
@@
The objective (22a) has two terms; the primary objective is the first term, which maximizes the fraction of visible targets at each time-step, and takes its summation. Note that (22a) is different from simply taking $\sum_{t=1}^{\ell}\sum_{k=1}^{q}\sum_{i=1}^{m}\sum_{j=1}^{n}M^{b}_{ijtk}X_{ijt}$, which would allow the optimizer to capitalize on observing the same target by multiple observers rather than observing as many targets as possible. The second term is a discount factor that penalizes the use of “expensive” observer locations, where $f\in\mathbb{R}^{n}$ is the vector of costs associated with making use of a given facility location, given by
@@
where $\nu_{j}$ is the stability index of the LPO, as defined by equation (14). The definition of $f_{j}$ with equation (23) allows the “cost” of a given facility location to have values close to one another: for the considered set of LPOs, $\nu_{j}$ varies between $1$ and $1399.19$, while $f_{j}$ varies between about $0.9091$ and $0.9993$. Since prohibitively expensive candidate observer locations can be excluded from the candidate observer location set $\mathcal{J}$ to begin with, it is assumed that all candidate LPOs have acceptable station-keeping cost should they be chosen as part of the constellation. As such, the second term in the objective is simply meant to favor combinations of facilities that are more stable in the case that multiple combinations of facilities yield the same value from the first term.
@@
where constraint (24b) ensures this modified problem (24) to have the same optimal solution as the original problem (22). Problem (24) is a standard mixed-binary linear program (MBLP), which is amenable to standard solvers such as Gurobi [65] or CPLEX [66]. While the slack variable $\theta_{tk}$ is a continuous variable in $[0,1]$, the optimal solution to (24) has $\theta_{tk}=1$ for time $t$ and target $k$ if at least one observer is viewing $k$ at this time, and $\theta_{tk}=0$ if $\sum_{i=1}^{m}\sum_{j=1}^{n}M^{b}_{ijtk}X_{ijt}=0$. Thus, this slack variable may be interpreted as a binary variable, such that
@@
For convenience in comparing obtained solutions later, we define the fraction of observation demand that has been met across all time steps, denoted by $\Theta$, and given by
@@
The formulation (24) enjoys a few favorable properties; firstly, even though the slack variables $\theta$ increase the decision space, it is a continuous variable, mending itself well to standard linear program techniques. Secondly, despite the introduction of the time index $t$ and the decomposition of allocation into pointing directions $i$ and targets $k$, the general structure of the traditional $p$-median problem is preserved; this enables efficient use of Lagrangian relaxation, which will be introduced in Section 5.
5 Lagrangian Method for the Time-Expanded $p$-Median Problems
@@
We develop a tailored version of the LM for the TE-$p$-MP. For traditional $p$-Median problems, the complicating constraints is typically those that pertain to allocations between $i$ and $j$, resulting in decoupled subproblems corresponding to each candidate facility location $j$. LM is particularly advantageous for the $p$-Median problem as the resulting relaxed subproblems can be solved analytically. Despite the introduction of the time indices $t$ and slack variables $\theta_{tk}$, we are also able to arrive at a relaxed problem that achieve decomposition into subproblems corresponding to each $j$. As a result, the upper bound solution $Z_{\rm relax}$ can be obtained analytically for a given set of Lagrange multipliers. We further provide a set of customized heuristics to generate a competitive feasible lower bound solution $Z_{\rm feas}$.
@@
The remainder of this Section is organized as follows: first, Section 5.1 discusses the relaxed problem formulation, along with the approach to decompose the resulting relaxed problem into facility-wise subproblems and to solve each of these subproblems. This is followed by Section 5.2 which describes the heuristic scheme devised to generate a feasible solution based on the relaxed solution at each iteration. An additional set of heuristics for performing a neighborhood search in an attempt to improve this feasible solution is described in Section 5.3. In Section 5.4, the subgradient method used to update the multipliers is introduced. The processes of the LM are summarized in Figure 8. Finally, in Section 5.5, we provide implementation considerations for implementing the proposed LM.
[…]
@@
Figure 8: Overview of Lagrangian method for the time-expanded $p$-Median problem
5.1 Upper Bound Solution via Lagrangian Relaxation
We first seek to find the decision variables and corresponding upper bound solution to the relaxed problem. Relaxing constraints (22b) and (24b) with multipliers $\lambda_{jt}\geq 0$ and $\eta_{tk}\geq 0$ respectively, the relaxed problem is given by
$\max_{X,Y,\theta}$ $\sum_{t=1}^{\ell}\sum_{k=1}^{q}\theta_{tk}-\dfrac{1}{\ell}\sum_{j=1}^{n}f_{j}Y_{j}-\sum_{j=1}^{n}\sum_{t=1}^{\ell}\lambda_{jt}\left(\sum_{i=1}^{m}X_{ijt}-1\right)-\sum_{t=1}^{\ell}\sum_{k=1}^{q}\eta_{tk}\left(\theta_{tk}-\sum_{i=1}^{m}\sum_{j=1}^{n}M^{b}_{ijtk}X_{ijt}\right)$ (27)
s.t. (22c) -(22e), (24c)
@@ 5.1 Upper Bound Solution via Lagrangian Relaxation
The sought optimal decision variables to the relaxed problem (27), denoted by $\bar{X}$, $\bar{Y}$, and $\bar{\theta}$ generates an upper bound $Z_{\rm relax}$ to problem (24).
@@ 5.1.1 Decomposition into Subproblems
We can further recognize that this relaxation decouples the problem on a per-candidate observer location basis. For each candidate location $j$, the subproblem is given by
@@
Subproblem (30) has as variables $X_{ijt}\in\mathbb{B}^{m\times 1\times\ell}$, $Y_{j}\in\mathbb{B}$, and $\theta\in\mathbb{R}^{\ell\times k}$. We denote the decision on $X$ and $Y_{j}$ made with the $j^{\rm th}$ subproblem as $\bar{X}^{j}$ and $\bar{Y}_{j}^{j}$, representing the allocations of the $j^{\rm th}$ observer to pointing directions over time, as well as whether the $j^{\rm th}$ observer should be used, respectively.
We consider separately the cases where $\bar{Y}^{j}_{j}=0$ or $\bar{Y}^{j}_{j}=1$; the decision on whether $\bar{Y}^{j}_{j}=0$ or $\bar{Y}^{j}_{j}=1$ is determined at a later stage, when the relaxed solution is constructed by assembling solutions from each subproblem. If $\bar{Y}^{j}_{j}=0$, constraint (30b) trivially forces $\bar{X}^{j}_{ijt}=0$ $\forall i,t$. If $\bar{Y}^{j}_{j}=1$, the choice of each $\bar{X}^{j}_{ijt}$ $\forall i,t$ is driven by its coefficient in the objective, and thus the optimal value $\bar{X}_{ijt}^{j}$ is given by
@@
The corresponding objective of the $j^{\mathrm{th}}$ subproblem (30a) assuming $\bar{Y}_{j}=1$, denoted by $Z^{j}_{\mathrm{relax}}$, is given by
$Z^{j}_{\mathrm{relax}}=\sum_{i=1}^{m}\sum_{t=1}^{\ell}\left(\sum_{k=1}^{q}\eta_{tk}M^{b}_{ijtk}-\lambda_{jt}\right)\bar{X}_{ijt}^{j}-\dfrac{f_{j}}{\ell}+\sum_{t=1}^{\ell}\lambda_{jt}$ (32)
5.1.2 Construction of Lagrangian Relaxed Solution
To construct the relaxed solution, we first compute $Z^{j}_{\mathrm{relax}}$ for all $j=1,\ldots,n$ subproblems (30) according to (32). We then define the permutation $\pi$ for $Z_{\rm relax}^{j}$ of the $j=1,\ldots,n$ candidate observer locations in descending order
$Z^{\pi(1)}_{\mathrm{relax}}\geq Z^{\pi(2)}_{\mathrm{relax}}\geq\ldots\geq Z^{\pi(m)}_{\mathrm{relax}}$ (33)
To enforce the $p$-median constraint (22c), variables $\bar{Y}$ are chosen such that the $p$ best facilities are chosen
@@ 5.1.2 Construction of Lagrangian Relaxed Solution
Finally, the upper bound objective of the Lagrangian relaxed problem (27) $Z_{\mathrm{relax}}$ is given by substituting the decision variables to equation (24a),
$Z_{\mathrm{relax}}=\sum_{t=1}^{\ell}\sum_{k=1}^{q}\left(1-\eta_{tk}\right)\bar{\theta}_{tk}+\sum_{s=1}^{p}Z^{\pi(s)}_{\mathrm{relax}}$ (36)
Algorithm 1 summarizes the relaxation process. Within the algorithm, argsortDescend returns the sorted indices of the inputs in descending order. In effect, the relaxed solution to the original problem (24) is obtained through only simple algebraic evaluations and without any iterative operation.
Algorithm 1 Lagrangian Relaxation for TE-$p$-MP
1:$\lambda_{jt}$ $\forall j,t$, $\eta_{tk}$ $\forall t,k$
2:$\bar{Y},\bar{X}\leftarrow\boldsymbol{0}_{n\times 1},\boldsymbol{0}_{m\times n\times\ell}$triangleright Initialize
[…]
4:for $j=1,\ldots,n$ do
[…]
7:end for
8:$\pi\leftarrow\verb|argsortDescend|(Z_{\rm relax}^{1},\ldots,Z_{\rm relax}^{n})$triangleright Sort in descending order according to (33)
9:for $j=\pi(1),\ldots,\pi(p)$ do
10:     $\bar{Y}_{j}\leftarrow 1$triangleright Use facility location $j$
11:     $\bar{X}_{ijt}\leftarrow\bar{X}_{ijt}^{j}$triangleright Assign targets according to decision made for subproblem $j$
12:end for
[…]
14:return $\bar{X}$, $\bar{Y}$, $\bar{\theta}$, $Z_{\rm relax}$
5.2 Lagrangian Heuristics Reallocation
If the Lagrangian relaxed solution $(\bar{X},\bar{Y},\bar{\theta})$ is feasible, then it is the optimal solution $(X^{*},Y^{*},\theta^{*})$ to problem (24); in most cases, $(\bar{X},\bar{Y},\bar{\theta})$ is infeasible as constraints (22b) and (24b) have been relaxed. Thus, we seek to construct a feasible lower bound solution $(\underline{X},\underline{Y},\underline{\theta})$ using the relaxed solution. The overall procedure is summarized in Algorithm 2.
@@ 5.2 Lagrangian Heuristics Reallocation
and we construct the set of all targets that are unobserved at time $t$, denoted by $\mathcal{K}^{0}_{t}$
@@
Equivalently, this set may be understood as the set of $j$ that still requires allocation to an appropriate pointing direction $i$. The next step consists of adding back allocations for $j\in\mathcal{J}_{t}^{0}$ for each time-step $t$, which corresponds to the GreedyAllocation() functions call in line 14 of Algorithm 2. We propose two approaches, coined the greedy allocation and full-factorial allocation, respectively. The choice between these two algorithms is made based on the number of observers $p$ of a given TE-$p$-MP instance and the relative computational time the user chooses to dedicate to the heuristic search. The two allocation algorithms are introduced in further detail in subsections 5.2.1 and 5.2.2. Finally, we fix the solution to satisfy constraint (24b) by iterating through $t$ and setting $\underline{\theta}_{tk}=0$ if
@@
Finally, the lower bound feasible solution $Z_{\mathrm{feas}}$ is given by
$Z_{\mathrm{feas}}=\sum_{t=1}^{\ell}\sum_{k=1}^{q}\left(1-\eta_{tk}\right)\underline{\theta}_{tk}+\sum_{j=1}^{n}f_{j}\underline{Y}_{j}$ (42)
@@ 5.2.1 Greedy Allocation
The first allocation strategy we propose is a greedy allocation. With the greedy allocation, the facility $j^{\star}$ with pointing direction $i^{\star}$ that results in the largest additional number of targets observed is selected via the expression
@@
This process is repeated, each time setting $X_{i^{\star}j^{\star}t}=1$ and removing $j^{\star}$ from $\mathcal{J}_{t}^{0}$, until $\mathcal{J}_{t}^{0}=\varnothing$. Pseudo-code for the greedy allocation scheme is provided in Algorithm 3. Since the algorithm essentially involves summation across $k$ and sorting across $i$ and $j$, its computational time is insignificant with dimensions of typical TE-$p$-MP instances.
@@
5.2.2 Full-Factorial Allocation
Typically, the number of observers $p$ in the context of the CSSA problem is relatively small, on the orders of 1s to 10s. Provided that $p$ is small enough, it may be computationally reasonable to spend additional time sequentially allocating remaining observers $j\in\mathcal{J}^{0}_{t}$ to remaining targets $\mathcal{K}^{0}_{t}$. We refer to this process as the full-factorial allocation; Algorithm 4 shows the pseudo-code for its procedure. The while-loop from the greedy Algorithm 3 is now replaced by the outer for-loop in line 2, which iterates through all possible permutations of indices of the set $\mathcal{J}^{0}_{t}$, denoted by $\Pi_{\mathcal{J}}$. For each permutation $\pi_{\mathcal{J}}\in\Pi_{\mathcal{J}}$, we go through indices $j$ in the order of $\pi_{\mathcal{J}}$, each time allocating $j$ to pointing direction $i$ in a greedy manner, such that